If three neurotransmittors bind to a normal neuron at resting Vm on the dendrites..

1st neurotransmittor: opens Na+ channels and causes graded potential of 5mV

2nd neurotransmittor: opens Cl- channels and causes graded potential of 8mV

3rd neurotransmittor: opens K+ channels and causes graded potential of 10mV

will this neuron have an action potential? please explain.
Answer #1

The resting membrane potential of a neuron is -70 mV. And the Threshold potential of a neuron is -55 mV.

So for a neuron to get depolarised we need a minimum of following potential:

Potential required = -55 -(-70) = 70-55 = +15mV

So we need a stimulus of total 15mV for the nerve to have an action potential.

We have stimulus of = 5 + 8+10 = +23 mV

This is greater than the Threshold potential required. So the Neuron will have the action potential.

It was a pleasure answering your question. Hope this helps...

Good luck!